OpenStudy (anonymous):

What is the value of u + v + w when u = –10, v = 4, and w = 8? A. –22 B. 2 C. 14 D. 22

TheSmartOne (thesmartone):

-10+4+8 =-6+8 =2

TheSmartOne (thesmartone):

@Smallathenabug

OpenStudy (anonymous):

@TheSmartOne thanks x